# Quantum Generative Diffusion Model: A Fully Quantum-Mechanical Model for Generating Quantum State Ensemble

### v2

@@ VI-C Mixed State Generation
TABLE III: Comparison of mixed state generation performance between QuGAN, EQ-GAN, and our proposed methods.
@@
In Table. III, we present a detailed comparison of the fidelity among these four methods. QuGAN starts at a fidelity of 0.902$\pm$6e-2 at $N=1$, declining to 0.324$\pm$6e-2 at $N=8$. This trend highlights QuGAN’s limitations regarding scalability. EQ-GAN has a lower standard deviation in most cases (except for the case $N=3$), while generally displaying higher average fidelity compared to QuGAN. In contrast, QGDM and RE-QGDM maintain a fidelity of greater than 0.99 across all results, accompanied by a low standard deviation, indicating their superior scalability. Notably, our methods achieve fidelity of up to 0.999 in some cases, demonstrating their excellent generative capabilities. In particular, RE-QGDM achieves 0.992$\pm$1e-2 for $N=7$ and maintains 0.993$\pm$1e-2 for $N=8$, demonstrating the ability of our models to maintain high fidelity generation even on larger scale problems.
@@
Our method demonstrated significant advantages in comparison with QuGAN and EQ-GAN. The results indicate that QuGAN and EQ-GAN perform poorly in the tasks of generating mixed states. The fidelity decreases as the number of qubits increases, revealing the limitations of these models in handling multi-qubit tasks. This limitation might be due to QGANs’ optimization strategies, which rely on the min-max saddle point game, which is generally less effective than using gradient descent for convex functions. It is well known that training classical GAN models is prone to issues such as model collapse, instability, and non-convergence. With an exponential increase in the dimensions of the density matrix, there may be higher demands on the training stage, such as the need for a more suitable learning rate. Given these conditions, the QGANs’ unstable training performance may significantly impact the final generation results. In contrast, the training process of (RE-) QGDM, which minimizes a convex function using gradient descent, typically shows more user-friendly and stable training performance. This might explain why our model outperforms QuGAN and EQ-GAN.
VII Models Training and Generation Curves
Figure 14: Metrics of training and generation process of QGDM.
Figure 15: Metrics of training and generation process of RE-QGDM when $N$ = 2, 3, 4 and 5.
Figure 16: Metrics of training and generation process of RE-QGDM when $N$ = 6, 7 and 8.
Fig. 14 displays the metrics during the training and generation of QGDM, which include the loss values and the fidelity of the generated quantum states. In contrast, Fig. 15 and Fig. 16 show the metrics of RE-QGDM.
Each scenario (number of qubits, pure or mixed state) includes two subfigures: the left subfigure illustrates the trend of training loss function value as the number of epochs increases, while the right subfigure shows how fidelity changes over timesteps. The red dashed line indicates the fidelity changes between the noisy quantum state and the target state during the diffusion process, as time steps increase (indicated by the red arrow from right to left). The solid blue line represents the trend of fidelity between the generated quantum state and the target state during the generation process (indicated by the blue arrow, from left to right). The shaded area in the figure represents the range of standard deviation.
The loss curves indicate that the loss functions of both QGDM and RE-QGDM converge to near zero as training epochs increase, and the number of epochs required to converge to zero increases with $N$, regardless of whether the target state is pure or mixed. Moreover, generating mixed states requires more iterations than pure states, especially as $N$ increases. This suggests that generating mixed states is more challenging than generating pure states.
As shown in Appendix B, the fidelity between any quantum state $\rho$ and the completely mixed state is $\frac{1}{d}\leq F(\rho,\mathbb{I}/d)\leq 1$. For pure states, the fidelity between the target state and the completely mixed state is $1/d$. Therefore, the standard deviation of all diffusion fidelity curves (red dashed lines) for pure states is zero. In the case of mixed states, the lower bound of the fidelity is $\frac{1}{d}$, and the variance decreases as $N$ increases. Therefore, at $t=T$ ($i.e.$, the far left of the horizontal axis), both the fidelity values and the standard deviation decrease with $N$.
In all scenarios, the fidelity curve during generation (blue solid line) increases as the timestep $t$ decreases, eventually approaching 1 with the smallest standard deviation. The overlap between the generation fidelity curve and the diffusion fidelity curve decreases as the number of qubits increases. As the number of qubits increases, the generation-fidelity curve diverges more from the diffusion-fidelity curve. Additionally, the generation fidelity curve (blue solid line) becomes steeper as $N$ increases.
VIII Conclusion
@@
Inspired by classical diffusion models, this paper introduces the Quantum Generative Diffusion Model (QGDM), a novel quantum generative approach for generating a quantum state ensemble. The fundamental concept of QGDM is that any target quantum state can be converted into a completely mixed state through a non-unitary forward process. Following this, a trainable backward process can be employed to reconstruct the target state from the completely mixed state. In the forward process, a timestep-dependent depolarizing channel incrementally converts the input state into a maximally random state, known as a completely mixed state. The denoising process involves a parameterized function composed of a timestep embedding circuit and a denoising circuit, employing the concept of parameter sharing. Specifically, the timestep embedding circuit embeds temporal information into a quantum state, while the denoising circuit processes the polluted quantum state along with the timestep embedding state to produce a denoised state. To minimize the number of auxiliary qubits in QGDM, we introduce a novel resource-efficient version of QGDM (RE-QGDM). Numerical simulations, conducted up to 8 qubits, compare our models with QuGAN and EQ-GAN. The results indicate that both QGDM and RE-QGDM surpass QuGAN and EQ-GAN in generating both pure and mixed quantum states, highlighting their excellent performance and scalability.
@@
Future work could explore the impact of different noise schedules on generation performance. Additionally, designing more efficient structures for the denoising process to reduce resource requirements, including the number of auxiliary qubits and the total number of quantum gates, is a valuable and worthwhile direction to pursue. Investigating the relationship between problem properties and the number of different time steps can provide important insights for further understanding QGDM. Theoretically analyzing and proving that QGDM has better convergence than previous QGAN models is an important direction for future exploration. This could provide theoretical support for a series of quantum versions of classic diffusion models and deepen our understanding of the advantages of these models.
@@ Appendix A The hyperparameters of the proposed models
TABLE I: The hyperparameters of QGDM.
@@
TABLE II: The hyperparameters of RE-QGDM.
@@
Appendix B Derivation of fidelity boundary between a quantum state and completely mixed state
@@
We start by considering the definition of fidelity $F$ between a quantum state $\rho$ and the completely mixed state $\frac{\mathbb{I}}{d}$, where $\mathbb{I}$ is the identity matrix and $d$ is the dimension of the Hilbert space. The fidelity is given by:
$F\left(\rho,\frac{\mathbb{I}}{d}\right)$ $=\left(\operatorname{tr}\sqrt{\sqrt{\rho}\frac{\mathbb{I}}{d}\sqrt{\rho}}\right)^{2}$ (A1)
@@
$=\frac{1}{d}\left(\sum_{k}\sqrt{\lambda_{k}}\right)^{2}$
@@ 1. If $\rho$ is the completely mixed state, then fidelity reaches its maximum:
$F\left(\rho,\frac{\mathbb{I}}{d}\right)=1.$ (A2)
@@
$F\left(\rho,\frac{\mathbb{I}}{d}\right)$ $=\frac{1}{d}\left(\sum_{k}\sqrt{\lambda_{k}}\right)^{2}$ (A3)
@@
$F\left(\rho,\frac{\mathbb{I}}{d}\right)=\frac{1}{d}$ (A4)
@@
$\frac{1}{d}\leq F\left(\rho,\frac{\mathbb{I}}{d}\right)\leq 1.$ (A5)

### v3

@@ 6.3 Mixed State Generation
TABLE III: Comparison of mixed state generation performance between QuGAN, EQ-GAN, and our proposed models.
@@
In Table III, we present a detailed comparison of the fidelity among these four methods. QuGAN starts at a fidelity of 0.902$\pm$6e-2 at $N=1$, declining to 0.324$\pm$6e-2 at $N=8$. This trend highlights QuGAN’s limitations regarding scalability. EQ-GAN has a lower standard deviation in most cases (except for the case $N=3$), while generally displaying higher average fidelity than QuGAN. In contrast, QGDM and RQGDM maintain a fidelity of greater than 0.99 across all cases, accompanied by a low standard deviation, indicating their superior scalability. Notably, our methods achieve fidelity of up to 0.999 in some cases, demonstrating their excellent generative capabilities. In particular, RQGDM achieves 0.992$\pm$1e-2 for $N=7$ and maintains 0.993$\pm$1e-2 for $N=8$, demonstrating the ability of our models to maintain high fidelity generation even on larger scale problems.
We can quantify the performance improvement of our models relative to the QGAN-based models by calculating the relative change [72] across different model groups ($i.e.$, our models and QGAN-based models). By extracting the best fidelity values for each $N$ from different model groups in Table III and calculating the average fidelities, we find that our models demonstrates a 53.02% relative performance improvement over QGAN-based models. For detailed calculations, please refer to the supplementary file.
@@
Our method has demonstrated significant advantages in comparison with QuGAN and EQ-GAN. The results indicate that QuGAN and EQ-GAN perform poorly in the tasks of generating mixed states. The fidelity decreases as the number of qubits increases, revealing the limitations of these models in handling multi-qubit tasks. This limitation might be due to QGANs’ optimization strategies, which rely on the min-max saddle point game, which is generally less effective than using gradient descent for convex functions. It is well known that training classical GAN models is prone to issues such as model collapse, instability, and non-convergence. With an exponential increase in the dimensions of the density matrix, there may be higher demands on the training stage, such as the need for a more suitable learning rate. Given these conditions, the QGANs’ unstable training performance may significantly impact the final generation results. In contrast, the training process of our models, which minimizes a convex function using gradient descent, shows more user-friendly and stable training performance. This might explain why our model well outperforms QuGAN and EQ-GAN.
7 Conclusion
@@
Inspired by classical diffusion models, this paper has introduced a Quantum Generative Diffusion Model (QGDM), as a novel quantum generative approach for generating a quantum state ensemble. Its fundamental concept is that any target quantum state can be converted into a completely mixed state through a non-unitary forward process. Following this, a trainable backward process is employed to reconstruct the target state from the completely mixed state. In the forward process, a timestep-dependent depolarizing channel incrementally converts the input state into a maximally random state, known as a completely mixed state. The denoising process involves a parameterized function composed of a timestep embedding circuit and a denoising circuit, employing the concept of parameter sharing. Specifically, the timestep embedding circuit embeds temporal information into a quantum state, while the denoising circuit processes the polluted quantum state along with the timestep embedding state to produce a denoised state. To minimize auxiliary qubits in QGDM, we introduce a novel resource-efficient version of QGDM called RQGDM. Numerical simulations, conducted up to 8 qubits, compare our models with QuGAN and EQ-GAN. The results indicate that both QGDM and RQGDM surpass QuGAN and EQ-GAN in generating both pure and mixed quantum states. Notably, our models achieve 53.02% higher fidelity in mixed-state generation tasks than QGANs. These results highlight their excellent performance and scalability.
Our work introduces new models to the field of quantum generative learning. These models overcome the convergence issues faced by QGANs when generating mixed states, providing powerful new tools for the challenging task of quantum state generation. Many of the designs and findings in our paper are pioneering and can serve as prototypes and valuable inspirations for future research on quantum versions of diffusion models.
@@
Future work should explore the impact of different noise schedules on generation performance. Additionally, designing more efficient structures for a denoising process to reduce resource requirements, including the number of auxiliary qubits and the total number of quantum gates, is a worthwhile direction to pursue. Investigating the relationship between problem properties and the number of different time steps can provide important insights for further understanding QGDM. Theoretically analyzing and proving that QGDM converges faster than QGAN is an important and open issue. This could provide theoretical support for quantum versions of classic diffusion models and deepen our understanding of the advantages of these models.
Acknowledgments
This work is supported by the Science and Technology Development Fund, Macau SAR (0093/2022/A2, 0076/2022/A2, and 0008/2022/AGJ), the National Key Research and Development Program of China (2023YFB2703800) and Guangdong Basic and Applied Basic Research Foundation (No. 2022A1515140116). The authors are thankful to Prof. Wenmin Wang for his valuable discussions.
Supplementary File for “Quantum Generative Diffusion Model: A Fully Quantum-Mechanical Model for Generating Quantum State Ensemble”
Chuangtao Chen, Qinglin Zhao, Senior Member, IEEE, MengChu Zhou, Fellow, IEEE, Zhimin He,
Zhili Sun, Senior Member, IEEE and Haozhen Situ
[…]
This supplementary file provides detailed additional information of our proposed models. Section A introduces the hyperparameters used in QGDM and RQGDM models. Section B details the calculation of relative performance improvement in mixed state generation. Section C presents the architecture details of the timestep embedding circuit and the denoising circuit. Section D presents the details of the target state preparation in the numerical simulations. Section E presents the training and generation curves for QGDM and RQGDM. Finally, Section F derives the fidelity boundary between a quantum state and the completely mixed state.
@@ Appendix A The Hyperparameters of the Proposed Models
TABLE IV: The hyperparameters of QGDM.
@@
TABLE V: The hyperparameters of RQGDM.
@@
Appendix B Calculation of Relative Performance Improvement in Mixed State Generation
We use the relative change [1] to quantify the improvement of our models over QGAN-based models in the mixed state generation task. First, we calculate the average of the best fidelity values for our models across different $N$:
$\bar{F}^{*}_{\text{Ours}}=\frac{0.995+0.999+0.999+0.996+0.993+0.990+0.992+0.993}{8}=0.994625.$ (14)
Then, we calculate the average of the best fidelity values for QGAN-based models across different $N$:
$\bar{F}^{*}_{\text{QGANs}}=\frac{0.972+0.867+0.741+0.676+0.602+0.555+0.463+0.324}{8}=0.65.$ (15)
The relative performance improvement can be calculated as follows:
$\frac{\bar{F}^{*}_{\text{Ours}}-\bar{F}^{*}_{\text{QGANs}}}{\bar{F}^{*}_{\text{QGANs}}}\times 100\%=\frac{0.994625-0.65}{0.65}\times 100\%=53.0192\%\approx 53.02\%.$ (16)
Appendix C Circuit Architecture
C.1 Timestep Embedding Circuit
The architecture of an timestep embedding circuit with $N_{\tau}=4$ qubits is shown in Fig. 11. The circuit includes single-qubit gates $Rx$ and $Ry$, and the two-qubit gate $ZZ(\phi)=\exp{(-i\phi(Z\otimes Z)/2)}$. Our proposed method maps temporal information to appropriate positions on the Bloch sphere (to be demonstrated by our numerical simulations) while maintaining linear complexity in the number of quantum gates, which is sufficient to meet our requirements.
Figure 11: The conventional ansatz [2] is used for our timestep embedding with trainable parameter $\boldsymbol{\omega}$, and given parameter $t^{\prime}=t\pi/T$. The circuit in the dashed lines is repeated for $L_{\mathcal{T}}$ times. Where $l\in\{1,\dots,L_{\mathcal{T}}\}$.
C.2 Denoising Circuit
The architecture of a 4-qubit denoising circuit $U(\boldsymbol{\theta})$ is shown in Fig. 12. The circuit includes single-qubit gates $Rz$ and $Rx$, and the two-qubit gate $XX(\phi)=\exp{(-i\phi(X\otimes X)/2)}$. This circuit architecture using $\mathcal{O}(N^{2})$ two-qubit gates may affect the efficiency of our method. However, our focus is to evaluate the effectiveness of our method. To ensure that this evaluation is not affected by the circuit design, we choose a more complex architecture, even though it increases the gate complexity. This decision allows our results to accurately represent the potential of our method. In addition, it may not be necessary for us to maintain a consistent circuit architecture across all timesteps. Adaptable circuits that change architecture at different timesteps can reduce the total number of gates in QGDM. We leave the task of designing efficient circuit architectures for QGDM for future work. An effective approach is to use quantum architecture search algorithms [3, 4, 5, 6].
Figure 12: The architecture of a 4-qubit denoising circuit $U(\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ denotes trainable parameters. To enhance its learning capability, the dashed-boxed part is repeated $L$ times. $l\in\{1,\dots,L\}$.
Appendix D Target State Preparation
D.1 Pure State
D.2 Mixed State
Quantum mixed states are generated through $\rho=\sum_{i}{p_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|}$, where $\{\left|\psi_{i}\right\rangle\}$ is randomly produced by the circuit shown in Fig. 13. The probability vector $\{p_{i}\}$ is obtained by uniform sampling from $(0,1]$ and normalized through a softmax operation to ensure $\sum_{i}{p_{i}}=1$ and $p_{i}>0$ for any index $i$. In this paper, we generate a target mixed state by probabilistically mixing two random pure states.
Appendix E Model Training and Generation Curves
Fig. 14 displays the metrics during the training and generation of QGDM, which include the loss values and fidelity of the generated quantum states. In contrast, Figs. 15-16 show the metrics of RQGDM.
Each scenario (number of qubits, pure or mixed state) includes two subfigures: the left subfigure illustrates the trend of training loss function value as the number of epochs increases, while the right subfigure shows how fidelity changes over timesteps. The red dashed line indicates the fidelity changes between the noisy quantum state and target one during the diffusion process, as time steps increase (indicated by the red arrow from right to left). The solid blue line represents the trend of fidelity between the generated quantum state and target one during the generation process (indicated by the blue arrow, from left to right). The shaded area in the figure represents the range of standard deviation.
The loss curves indicate that the loss functions of both QGDM and RQGDM converge to near zero as training epochs increase, and the number of epochs required to converge to zero increases with $N$, regardless of whether the target state is pure or mixed. Moreover, generating mixed states requires more iterations than generating pure ones, especially as $N$ increases. This suggests that generating mixed states is more challenging than generating pure ones.
As shown in Section F, the fidelity between any quantum state $\rho$ and the completely mixed state is $\frac{1}{d}\leq F(\rho,\mathbb{I}/d)\leq 1$. For pure states, the fidelity between the target state and the completely mixed state is $1/d$. Therefore, the standard deviation of all diffusion fidelity curves (red dashed lines) for pure states is zero. For mixed states, the lower bound of the fidelity is $\frac{1}{d}$. As $N$ increases, the random quantum states appear more similar [7], causing the variance of the fidelity to decrease. Therefore, at $t=T$ ($i.e.$, the far left of the horizontal axis), both the fidelity values and the standard deviation decrease with $N$.
In all scenarios, the fidelity curve during generation (blue solid line) increases as the timestep $t$ decreases, eventually approaching 1 with the smallest standard deviation. The overlap between the generation fidelity curve and the diffusion one decreases as the number of qubits increases. As the number of qubits increases, the generation-fidelity curve diverges more from the diffusion-fidelity curve. Additionally, the generation fidelity curve (blue solid line) becomes steeper as $N$ increases.
Figure 14: Metrics of training and generation process of QGDM.
Figure 15: Metrics of training and generation process of RQGDM when $N$ = 2, 3, 4 and 5.
Figure 16: Metrics of training and generation process of RQGDM when $N$ = 6, 7 and 8.
@@ Appendix E Model Training and Generation Curves
Appendix F Derivation of Fidelity Boundary between a Quantum State and Completely Mixed State
@@
We start by considering the definition of quantum fidelity $F$ between a quantum state $\rho$ and the completely mixed state $\frac{\mathbb{I}}{d}$, where $\mathbb{I}$ is the identity matrix and $d$ is the dimension of the Hilbert space. The fidelity is given by:
$F\left(\rho,\frac{\mathbb{I}}{d}\right)$ $=\left(\operatorname{tr}\sqrt{\sqrt{\rho}\frac{\mathbb{I}}{d}\sqrt{\rho}}\right)^{2}$ (17)
@@
$=\frac{1}{d}\left(\sum_{k}\sqrt{\lambda_{k}}\right)^{2}.$
@@ 1. If $\rho$ is the completely mixed state, then fidelity reaches its maximum:
$F\left(\rho,\frac{\mathbb{I}}{d}\right)=1.$ (18)
@@
$F\left(\rho,\frac{\mathbb{I}}{d}\right)$ $=\frac{1}{d}\left(\sum_{k}\sqrt{\lambda_{k}}\right)^{2}$ (19)
@@
$F\left(\rho,\frac{\mathbb{I}}{d}\right)=\frac{1}{d}.$ (20)
@@
$\frac{1}{d}\leq F\left(\rho,\frac{\mathbb{I}}{d}\right)\leq 1.$ (21)